zepdrix (zepdrix):

Cindee, maf time! @Kayders1997

OpenStudy (kayders1997):

Yes :(

zepdrix (zepdrix):

question? :D

OpenStudy (kayders1997):

Just 1 sec

OpenStudy (kayders1997):

I was helping someone with a question before you came we were almost done

zepdrix (zepdrix):

k

OpenStudy (kayders1997):

Hopefully lol

OpenStudy (kayders1997):

It's a basic equation 7x-1=20

OpenStudy (kayders1997):

Okay the integral Sinpitdt

zepdrix (zepdrix):

\[\large\rm \int\limits \sin(\pi t)~dt\]

OpenStudy (kayders1997):

How do you do that?

zepdrix (zepdrix):

Magic. Do you understand how to do this integral?\[\large\rm \int\limits \sin(u)~du\]

OpenStudy (kayders1997):

I barely understand u sub 😂

zepdrix (zepdrix):

Well we're not dealing with that just yet. Do you know how to find the anti-derivative of sin x? That's all im asking so far.

OpenStudy (kayders1997):

Oh is it -cosx?

zepdrix (zepdrix):

Ok great. So if we were somehow able to transform our integral to sin(u), we would know it turns into -cos(u).

OpenStudy (kayders1997):

Okay

zepdrix (zepdrix):

\[\large\rm \int\limits\limits \sin(\color{orangered}{\pi t})~\color{royalblue}{dt}\]So we'll make the substitution \(\large\rm \color{orangered}{u=\pi t}\), that will `partially` give us what we want,\[\large\rm \int\limits\limits\limits \sin(\color{orangered}{\pi t})~\color{royalblue}{dt}\quad=\quad \int\limits\limits\limits \sin(\color{orangered}{u})~\color{royalblue}{dt}\]

zepdrix (zepdrix):

If we're going to do that, we also must replace the differential dt with something involving du.

OpenStudy (kayders1997):

I don't understand what your supposed to choose at u, the other part that we didn't antiderivative at the beginning?

zepdrix (zepdrix):

In general, you're looking for a \(\large\rm u\) and \(\large\rm u'\), something and it's derivative. It's a little more simple than that in this case though. We have sin(pi t) but we know how to integrate sin(x) or sin(u), so we're replacing the (pi t) by u.

zepdrix (zepdrix):

We're using our understanding of the basic integral to simplify it down, that's how we know what to choose, whatever u makes the problem easier.

zepdrix (zepdrix):

It won't always be immediately obvious though.

OpenStudy (kayders1997):

Okay I'm understanding more now

zepdrix (zepdrix):

We also need to replace the differential with something involving u. So start here with your substitution,\[\large\rm u=\pi t\]and take a derivative, with respect to t.

OpenStudy (kayders1997):

Okay so du/dt=pi?

zepdrix (zepdrix):

Ok great. Here is another way we can write this,\[\large\rm \frac{du}{dt}=\pi\qquad\qquad\to\qquad\qquad du=\pi~dt\]You can think of this process as "multiplying" the dt to the other side. That's not entirely accurate since dt is an infinitesimal quantity, but it follows the same rule.

OpenStudy (kayders1997):

Okay that makes sense

OpenStudy (kayders1997):

Yeah I was like ugghhhhhh idk lol

zepdrix (zepdrix):

Notice that this dt is the same dt in our integral,\[\large\rm du=\pi~\color{royalblue}{dt}\]So think of it as a veriable that you would like to isolate. I guess we can divide the pi to the other side, ya?

OpenStudy (kayders1997):

Yeah

OpenStudy (kayders1997):

This is getting really messy

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{\frac{1}{\pi}du=dt}\]

zepdrix (zepdrix):

Eh it's not bad :D

zepdrix (zepdrix):

Let's group up our information so we don't lose anything. One sec,

OpenStudy (kayders1997):

Okay, and none relevant question is this the last harder concept in calculus?

OpenStudy (kayders1997):

Hardest

zepdrix (zepdrix):

No, power series is much much harder (at least for me). But I tend to think of these types of integrals as really easy, so maybe it's just different for different people.

OpenStudy (kayders1997):

Omg :O I don't like these but someone new has been teaching for about 1 week now

zepdrix (zepdrix):

We start with this integral,\[\large\rm \int\limits\limits\limits \sin(\color{orangered}{\pi t})~\color{royalblue}{dt}\]and we determined that this substitution would be appropriate: \[\large\rm \color{orangered}{u=\pi t}\]\[\large\rm \color{royalblue}{\frac{1}{\pi}du=dt}\]Making these replacements gives us,\[\large\rm \int\limits\limits\limits\limits \sin(\color{orangered}{u})~\color{royalblue}{\frac{1}{\pi}du}\]

zepdrix (zepdrix):

I hope the color is helping a little bit. Pay attention to that. I'm just replacing orange with orange, and blue with blue.

OpenStudy (kayders1997):

Yes the color helps a lot thank you for that

zepdrix (zepdrix):

From this point, let's pull the 1/pi out in front of the integral,\[\large\rm \frac{1}{\pi}\int\limits \sin(u)~du\]

OpenStudy (kayders1997):

Because it's a constant!

zepdrix (zepdrix):

yes, constant coefficient.

zepdrix (zepdrix):

And now, integrate.

OpenStudy (kayders1997):

So -cosu du/dt?

zepdrix (zepdrix):

0_o

OpenStudy (kayders1997):

Don't know about the second part :O

zepdrix (zepdrix):

what second part?

OpenStudy (kayders1997):

Well the du/dt

zepdrix (zepdrix):

there is no du/dt in the integral, i'm not sure what you're talking about :U

OpenStudy (kayders1997):

Idk either I just got confused when you wanted me to anti differentiate

zepdrix (zepdrix):

Well you told me earlier that the anti-derivative of sin(x) is -cos(x), right? So,\[\large\rm \frac{1}{\pi}\color{purple}{\int\limits\limits \sin(u)~du}\qquad=\quad \frac{1}{\pi}\color{purple}{(-\cos(u))}\]

OpenStudy (kayders1997):

What happened to du 0_o

zepdrix (zepdrix):

The du and the big squiggly bar are buddies. They come and go as a package. When you perform the "integration" step, they both disappear. Not just the squiggly bar, but also the differential.

OpenStudy (kayders1997):

Oh weird

OpenStudy (kayders1997):

So is that the answer?

zepdrix (zepdrix):

\[\large\rm \int\limits x~dx=\frac12x^2\]Both the bar and the dx disspear

zepdrix (zepdrix):

example ^

zepdrix (zepdrix):

Well we have one final step,\[\large\rm -\frac{1}{\pi}\cos(\color{orangered}{u})\]Remember that the problem was given to us in terms of \(\large\rm t\), not this silly \(\large\rm u\), so finally, we need to `undo our subistution`.

OpenStudy (kayders1997):

Ohhhhhhhh

OpenStudy (kayders1997):

So -1/pi cos(pi(t))?

zepdrix (zepdrix):

\[\large\rm \int\limits \sin(\pi t)~dt\quad=\quad -\frac{1}{\pi}\cos(\pi t)+C\]Good good good.

OpenStudy (kayders1997):

Oh yeah that darn +c

zepdrix (zepdrix):

There is an easier, more intuitive way to do these, but you're just starting to learn u-sub, so we shouldn't try to stuff too much into your brain right now.

OpenStudy (kayders1997):

Okay plus if I SOMEHOW lol learned it I would probably want to do your way when it says on an exam solve by u substitution

zepdrix (zepdrix):

lol ya that'd be a problem XD

OpenStudy (kayders1997):

Are you tired or could we do one of the other problems the parts

zepdrix (zepdrix):

Ooo parts? bring it.

OpenStudy (kayders1997):

Okay so intergral of tsin2tdt

OpenStudy (kayders1997):

I know you need to pick out a u v dv and du

zepdrix (zepdrix):

You need to pick out \(\large\rm u\) and \(\large\rm dv\). From there, you'll `differentiate` to get a \(\large\rm du\), and you'll `integrate` to get a \(\large\rm u\). So realize...

OpenStudy (kayders1997):

So realize?

zepdrix (zepdrix):

that whatever you pick for your \(\large\rm u\) is going to be differentiated. Try to think about the differentiation process... Generally speaking, it makes things simpler, it breaks them down. At least this is true with polynomials right, if we have t^2, it turns into 2t, then 2, then 0 through multiple differentiations.

OpenStudy (kayders1997):

Right

OpenStudy (kayders1997):

So the easiest thing to differeniate here is t

zepdrix (zepdrix):

But sine does work so nicely, sine, cosine, -sine, -cosine, sine it's very circular, not nice for differentiation. Doesn't break down like we want.

zepdrix (zepdrix):

Good good good, so the \(\large\rm t\) is what we want for our \(\large\rm u\)

zepdrix (zepdrix):

Which means everything else is our \(\large\rm dv\).

zepdrix (zepdrix):

\[\large\rm \int\limits t \sin(2t)dt\]Oh they gave us a stinker... I guess we do have to learn this shortcut trick :( You really really don't want to apply a u-substitution within a parts problem. It get's really burdensome.

OpenStudy (kayders1997):

Okay, yeah they never taught us really

zepdrix (zepdrix):

Do you remember how to take this derivative?\[\large\rm \left(e^{2x}\right)'\]

OpenStudy (kayders1997):

I think so :)

OpenStudy (kayders1997):

Isn't it... 2e^2x?

zepdrix (zepdrix):

Good, you end up `multiplying` by that extra 2, because of the chain rule.

zepdrix (zepdrix):

So what would happen here?\[\large\rm \int\limits 2e^{2x}~dx\]

OpenStudy (kayders1997):

Why did we change it to e though?

zepdrix (zepdrix):

We'll relate it to the original problem in a moment :U You have to learn this trick though!

OpenStudy (kayders1997):

Okay!

OpenStudy (kayders1997):

Ummm I'm not sure :(

zepdrix (zepdrix):

Well, if the derivative of e^(2x) is 2e^(2x), then the anti-derivative of 2e^(2x) is e^(2x) right?

OpenStudy (kayders1997):

Lol omg yes

zepdrix (zepdrix):

\[\large\rm \int\limits\limits 2e^{2x}~dx\quad=\quad e^{2x}\]

zepdrix (zepdrix):

But what happened to the 2 in this process? Try to think about mathematically... which simple operation got rid of the 2?

OpenStudy (kayders1997):

2/2 o_O

zepdrix (zepdrix):

\[\large\rm \int\limits\limits\limits 2e^{2x}~dx\quad=\quad \frac22e^{2x}\quad=\quad e^{2x}\]Good good good, we actually had to `divide` by the 2 when we integrated, right?

OpenStudy (kayders1997):

Yes

zepdrix (zepdrix):

So what about this integral?\[\large\rm \int\limits e^{2x}dx\]Any ideas? :)